Please identify the operations performed on each side of the below equation.
3(x – 1) + 4(2x + 1) = 11x + 1
How can you make this equation into a function or function(s)?
This is what I come up with, but not sure if it is correct.
3 ( x - 1) + 4 ( 2x + 1) = 11x +1 Original Problem
(3x - 3) + (8x + 4) = 11x + 1 Distributive
3x - 3 + 8x + 4 = 11x + 1 Associative
11x + 1 = 11x + 1 Commutative
11x + 1 = 11x + 1 Subtraction to cancel the 1's out
11x = 11x Division to find x
11x/11
Answer: x = 1

Answer by Alan3354(69443)

Please identify the operations performed on each side of the below equation.
3(x – 1) + 4(2x + 1) = 11x + 1
How can you make this equation into a function or function(s)?
This is what I come up with, but not sure if it is correct.
3 ( x - 1) + 4 ( 2x + 1) = 11x +1 Original Problem
(3x - 3) + (8x + 4) = 11x + 1 Distributive
3x - 3 + 8x + 4 = 11x + 1 Associative
11x + 1 = 11x + 1 Commutative
11x + 1 = 11x + 1 Subtraction to cancel the 1's out
11x = 11x Division to find x
------------------- Ok above here
11x/11 What happened to the equal sign???
Answer: x = 1
---------------
x = x is all you can say.
3(x – 1) + 4(2x + 1) = 11x + 1 is TRUE for all values of x.